# Filtering Spin and Orbital Moment in Centrosymmetric Systems

### v2

@@ IV Conclusions
A direct application of our prediction is provided by the use of vanadocene or similar types of molecules [36]. Indeed, vanadocene possesses $D_{5d}$ point group symmetry, which means that it does not have a horizontal mirror plane and features five vertical mirror planes associated with its planar pentagonal structure. Hence, when it is put in contact with a cubic system, as for the case of elemental Ag, there can be no more that one common vertical mirror depending on the relative orientation of the symmetry axes. Then, the electronic states of the vanadocene will act as a source of orbital hybridization for the Ag states with different mirror parity, thus leading to couplings of the type employed in the examined model. Additionally, our investigation of the role of atomic SOC reveals that, when examining electrical transport with metallic leads connected to the transmission medium, it is not essential for the leads to consist of heavy elements with substantial SOC [31, 37, 38]. The capability for spin and orbital filtering does not depend on SOC in the areas where charge current is injected or collected, consistently with recent experimental observations [39].
@@ Appendix A Derivation of the effective model with crystalline arrangement of the transmission line domains
In this section we provide a connection between the examined model in the main text and a physical setup that aligns with a specific geometry or crystalline arrangement of the areas constituting the transmission line. To this aim, we consider the transmission problem for an electron propagating through three square-lattice regions: I, II, and III, each with the same lattice constant $a$ (see Fig. 5). We assume there is only one atom per unit cell in each region; in addition, the electronic transmission takes place along the $x$-direction selected with a given reference frame. The crystallographic axes of the outer lattice regions (I and III) are aligned with the reference frame, while the ones of the intermediate region (II) are rotated by an angle $\theta$, as depicted in Fig. 5. By adopting a tight-binding approach, we consider a multi-orbital system whose electronic states at a given energy $\mu$ are expressed in terms of the atomic orbitals $|p_{z}\rangle$,$|p_{y}\rangle$, and $|p_{x}\rangle$ that span the $L=1$ manifold. All the energies are given in unit of $\epsilon_{0}=\hbar^{2}k_{0}^{2}/2m$, with $m$ being the mass of the electron. From here on, $\hbar=1$ and $12m=$ 1 for convenience and clarity.
@@
^$\hat{\mathcal{H}}^{\text{II}}\left(\mathbf{k}\right)$ $=\cos\left[\left(k_{x}\cos\theta-k_{y}\sin\theta\right)a\right]\left(\sum_{i}\gamma_{i}^{\text{II}}\hat{L}_{i}^{2}+g_{1}\hat{\chi}_{z}\right)+$
$+$ $\cos\left[\left(k_{x}\sin\theta+k_{y}\cos\theta\right)a\right]\left(\sum_{i}\delta_{i}^{\text{II}}\hat{L}_{i}^{2}+g_{2}\hat{\chi}_{z}\right)+$
$+$ $\sum_{i}\xi_{i}^{\text{II}}\hat{L}_{i}^{2}+g_{3}\hat{\chi}_{z},$ (16)
with $\hat{\chi}_{z}=\hat{L}_{x}\hat{L}_{y}+\hat{L}_{y}\hat{L}_{x}$. Here, the terms proportional to $\hat{L}_{i}^{2}$ reproduce symmetry allowed crystal-field effects; on the contrary, $\hat{\chi}_{z}$ preserves inversion ^$\hat{\mathcal{I}}:\left(x,y,z\right)\rightarrow\left(-x,-y,-z\right)$, while breaks ^$\hat{\mathcal{M}}_{x,y}$ and $\hat{\mathcal{C}}_{2x,2y}$ by coupling the electronic orbital degrees of freedom. The involved parameters ($\gamma$, $\delta$, $\xi$, and $g$) depend on the angle $\theta$, the on-site energies, and the two-center bond integrals between two $p$-orbitals, $V_{pp\sigma}$ and $V_{pp\pi}$, where $\sigma$ and $\pi$ denote the type of the chemical bond [41]. Then, assuming that the electron propagates only along the $x$-axis, and in the thin limit regime, i.e., with a transverse dimension ($d_{y}$) that is much smaller than the longitudinal one ($d_{x}$), one can replace the $k_{y}$ component with averages that fulfill the following conditions: $\left\langle k_{y}^{2n+1}\right\rangle=0$ and $\left\langle k_{y}^{2n}\right\rangle\neq 0$ with $n=0,1,\ldots\,$. The resulting effective Hamiltonian is
{^$\begin{cases}\hat{\mathcal{H}}^{\text{I}}=k_{x}^{2}\left(\sum_{i}a_{i}^{\text{I}}\hat{L}_{i}^{2}\right)+\sum_{i}\Delta_{i}^{\text{I}}\hat{L}_{i}^{2}\\ \hat{\mathcal{H}}^{\text{II}}=k_{x}^{2}\left(\sum_{i}a_{i}^{\text{II}}\hat{L}_{i}^{2}+g_{z}\hat{\chi}_{z}\right)+\sum_{i}\Delta_{i}^{\text{II}}\hat{L}_{i}^{2}+g_{0z}\hat{\chi}_{z}\\ \hat{\mathcal{H}}^{\text{III}}=k_{x}^{2}\left(\sum_{i}a_{i}^{\text{III}}\hat{L}_{i}^{2}\right)+\sum_{i}\Delta_{i}^{\text{III}}\hat{L}_{i}^{2}\end{cases},$ (17)
@@
$\begin{cases}a_{i}^{\text{II}}&=-\frac{a^{2}}{2}\left(\alpha_{y}\gamma_{i}^{\text{II}}\cos^{2}\theta+\beta_{y}\delta_{i}^{\text{II}}\sin^{2}\theta\right)\\ g_{z}&=-\frac{a^{2}}{2}\left(\alpha_{y}g_{1}\cos^{2}\theta+\beta_{y}g_{2}\sin^{2}\theta\right)\\ \Delta_{i}^{\text{II}}&=\alpha_{y}\gamma_{i}^{\text{II}}+\beta_{y}\delta_{i}^{\text{II}}+\xi_{i}^{\text{II}}\\ g_{0z}&=\alpha_{y}g_{1}+\beta_{y}g_{2}+g_{3}\end{cases},$ (18)
$\alpha_{y}\sim 1-\frac{1}{2}\left\langle k_{y}^{2}\right\rangle a^{2}\sin^{2}\theta$, and $\beta_{y}\sim 1-\frac{1}{2}\left\langle k_{y}^{2}\right\rangle a^{2}\cos^{2}\theta$. Here, we have also taken the long-wavelength approximation such as $\cos x\sim 1-x^{2}/2$.
@@
This approach can also be extended to the three-dimensional case with different lattice geometries. Therefore, by considering a generic heterostructure, we adopt a generalization of the aforementioned effective model, assuming that the electron propagates along the $z$-axis of a coordinate system $Oxyz$. To include the electronic spin degree, we employ the Pauli matrices: $\hat{S}_{i=x,y,z}=\frac{1}{2}\hat{\sigma}_{i=x,y,z}$. For the analysis of the transmission problem, the electron is prepared in a given state $\left|\Psi_{in}^{\text{I}}\right\rangle$ in region I. There, the Hamiltonian is expressed as ^$\hat{\mathcal{H}}^{\text{I}}\left(z\right)=\left[\hat{p}_{z}^{2}\left(\sum_{i}a_{i}^{\text{I}}\hat{L}_{i}^{2}\right)+\sum_{i}\Delta_{i}^{\text{I}}\hat{L}_{i}^{2}\right]\hat{\sigma}_{0}$, where $\hat{p}_{z}$ is the electronic linear momentum operator along the $z$-axis, and $\hat{\sigma}_{0}=\hat{\sigma}_{i}^{2}$. The electron is thus injected from region I into region II, which spans from $z=-d/2$ to $z=d/2$. In this region, the single-electron Hamiltonian is given by
@@
$\quad+\left(\sum_{i}\Delta_{i}^{\text{II}}\hat{L}_{i}^{2}+\bm{g_{0}}\cdot\hat{\bm{\chi}}\right)\hat{\sigma}_{0}+\lambda\hat{\bm{L}}\cdot\hat{\bm{S}}.$ (19)
with $\bm{g}=\left(g_{x},g_{y},g_{z}\right)$, $\bm{g_{0}}=\left(g_{0x},g_{0y},g_{0z}\right)$, $\hat{\bm{\chi}}=\left(\left\{\hat{L}_{y},\hat{L}_{z}\right\},\left\{\hat{L}_{x},\hat{L}_{z}\right\},\left\{\hat{L}_{x},\hat{L}_{y}\right\}\right)$, and $\lambda$ as the atomic spin-orbit coupling magnitude (in units of $\epsilon_{0}$). Finally, the electron can reach the last region of space, region III, where its orbital and spin angular momenta are measured, and the Hamiltonian can be expressed as ^$\hat{\mathcal{H}}^{\text{III}}\left(z\right)=\left[\hat{p}_{z}^{2}\left(\sum_{i}a_{i}^{\text{III}}\hat{L}_{i}^{2}\right)+\sum_{i}\Delta_{i}^{\text{III}}\hat{L}_{i}^{2}\right]\hat{\sigma}_{0}$. Thus, we can define the following overall Hamiltonian ^=^$\hat{\mathcal{H}}=\hat{\mathcal{H}}\left(z\right)$:

### v3

@@ IV Conclusions
A direct application of our prediction is provided by the use of vanadocene or similar types of molecules [36]. Indeed, vanadocene possesses $D_{5d}$ point group symmetry, which means that it does not have a horizontal mirror plane and features five vertical mirror planes associated with its planar pentagonal structure. Hence, when it is put in contact with a cubic system, as for the case of elemental Ag, there can be no more than one common vertical mirror depending on the relative orientation of the symmetry axes. Then, the electronic states of the vanadocene will act as a source of orbital hybridization for the Ag states with different mirror parity, thus leading to couplings of the type employed in the examined model. Additionally, our investigation of the role of atomic SOC reveals that, when examining electrical transport with metallic leads connected to the transmission medium, it is not essential for the leads to consist of heavy elements with substantial SOC [31, 37, 38]. The capability for spin and orbital filtering does not depend on SOC in the areas where charge current is injected or collected, consistently with recent experimental observations [39].
@@ Appendix A Derivation of the effective model with crystalline arrangement of the transmission line domains
In this section, we provide a connection between the examined model in the main text and a physical setup that aligns with a specific geometry or crystalline arrangement of the areas constituting the transmission line. To this aim, we consider the transmission problem for an electron propagating through three square-lattice regions: I, II, and III, each with the same lattice constant $a$ (see Fig. 5). We assume there is only one atom per unit cell in each region. In addition, within a given reference frame, we consider the system to be strongly confined along the $y$-direction, $d_{y}\ll d_{x}$, such that the electronic transmission takes place along the $x$-direction. The crystallographic axes of the outer lattice regions (I and III) are aligned with the reference frame, while the ones of the intermediate region (II) are rotated by an angle $\theta$, as depicted in Fig. 5. By adopting a tight-binding approach, we consider a multi-orbital system whose electronic states at a given energy $\mu$ are expressed in terms of the atomic orbitals $|p_{z}\rangle$, $|p_{y}\rangle$, and $|p_{x}\rangle$ that span the $L=1$ manifold. All the energies are given in unit of $\epsilon_{0}=\hbar^{2}k_{0}^{2}/2m$, with $m$ being the mass of the electron. From here on, $\hbar=1$ and $12m=$ 1 for convenience and clarity.
@@
^$\hat{\mathcal{H}}^{\text{II}}\left(\mathbf{k}\right)=\mathcal{\hat{T}}_{0}+2\cos\left(\mathbf{k}\cdot\bm{a}_{1}\right)\mathcal{\hat{T}}_{1}+2\cos\left(\mathbf{k}\cdot\bm{a}_{2}\right)\mathcal{\hat{T}}_{2},$ (16)
@@
$\begin{cases}\mathcal{\hat{T}}_{0}=\xi_{x}^{\text{II}}\hat{L}_{x}^{2}+\xi_{y}^{\text{II}}\hat{L}_{y}^{2}+\xi_{z}^{\text{II}}\hat{L}_{z}^{2}\\ \mathcal{\hat{T}}_{1}=\gamma_{x}^{\text{II}}\hat{L}_{x}^{2}+\gamma_{y}^{\text{II}}\hat{L}_{y}^{2}+\gamma_{z}^{\text{II}}\hat{L}_{z}^{2}+g_{1}\hat{\chi}_{z}\\ \mathcal{\hat{T}}_{2}=\delta_{x}^{\text{II}}\hat{L}_{x}^{2}+\delta_{y}^{\text{II}}\hat{L}_{y}^{2}+\delta_{z}^{\text{II}}\hat{L}_{z}^{2}+g_{2}\hat{\chi}_{z}\end{cases}$ (17)
with $\hat{\chi}_{z}=\hat{L}_{x}\hat{L}_{y}+\hat{L}_{y}\hat{L}_{x}$. Here, the terms proportional to $\hat{L}_{i}^{2}$ reproduce symmetry allowed crystal-field effects; on the contrary, $\hat{\chi}_{z}$ preserves inversion ^$\hat{\mathcal{I}}:\left(x,y,z\right)\rightarrow\left(-x,-y,-z\right)$, while breaks ^$\hat{\mathcal{M}}_{x,y}$ and $\hat{\mathcal{C}}_{2x,2y}$ by coupling the electronic orbital degrees of freedom. The involved parameters ($\gamma$, $\delta$, $\xi$, and $g$) depend on the angle $\theta$, the on-site energies, and the two-center bond integrals between two $p$-orbitals, $V_{pp\sigma}$ and $V_{pp\pi}$, where $\sigma$ and $\pi$ denote the type of the chemical bond [41]. Particularly, $g_{1,2}=0$ if $\theta=0$. In the long-wavelength approximation, the trigonometric functions can be expanded as $\cos x\sim 1-x^{2}/2$ and $\sin x\sim x$. Thus, by retaining only terms quadratic in momentum, the effective Hamiltonian in the continuous limit takes the form
^$\hat{\mathcal{H}}^{\text{II}}=\hat{A}\hat{p}_{x}^{2}+\hat{B}\hat{p}_{y}^{2}+\hat{C}\hat{p}_{x}\hat{p}_{y}+\hat{D},$ (18)
where $\hat{p}_{i}$ is the electronic linear momentum operator along the $i$-axis, and
$\begin{cases}\hat{A}=-a^{2}\cos^{2}\theta\mathcal{\hat{T}}_{1}-a^{2}\sin^{2}\theta\mathcal{\hat{T}}_{2}\\ \hat{B}=-a^{2}\sin^{2}\theta\mathcal{\hat{T}}_{1}-a^{2}\cos^{2}\theta\mathcal{\hat{T}}_{2}\\ \hat{C}=2a^{2}\sin\theta\cos\theta\left(\mathcal{\hat{T}}_{1}-\mathcal{\hat{T}}_{2}\right)\\ \hat{D}=\mathcal{\hat{T}}_{0}+2\mathcal{\hat{T}}_{1}+2\mathcal{\hat{T}}_{2}\end{cases}.$ (19)
We choose a complete basis composed of factorized states, $\left\{\left|\Psi_{n}\left(x,y\right)\right\rangle=\psi_{n_{1}}\left(x\right)\phi_{n_{2}}\left(y\right)\left|\eta_{n_{3}}\right\rangle\right\}$, such that $\left\langle\psi_{n_{1}}|\psi_{m_{1}}\right\rangle=\delta_{n_{1},m_{1}}$, $\left\langle\phi_{n_{2}}|\phi_{m_{2}}\right\rangle=\delta_{n_{2},m_{2}}$, and $\left\langle\eta_{n_{3}}|\eta_{m_{3}}\right\rangle=\delta_{n_{3},m_{3}}$, with $|\eta_{n_{3}}\rangle$ being the basis associated with the orbital degrees of freedom. Given the confinement of the system along the $y$-direction, we impose hard-wall boundary conditions along this axis: $\left|\Psi\left(x,y=\pm d_{y}/2\right)\right\rangle=0$. This leads to considering
$\phi_{n}^{+}\left(y\right)=d_{y}^{-1/2}\cos\left(k_{y,n}^{+}y\right),\,\,\,\phi_{n}^{-}\left(y\right)=d_{y}^{-1/2}\sin\left(k_{y,n}^{-}y\right)$ (20)
as electronic transverse modes, where $+$ (--) denotes even (odd) parity and
$k_{y,n}^{+}=\frac{\left(2n-1\right)\pi}{d_{y}},\,\,\,k_{y,n}^{-}=\frac{2n\pi}{d_{y}}$ (21)
with $n=1,2,\ldots$. Applying the operator $\hat{p}_{y}^{2}$ on the transverse modes yields the corresponding confinement energies
$E_{y,n}^{+}=\frac{\left(2n-1\right)^{2}\pi^{2}}{d_{y}^{2}},\,\,\,E_{y,n}^{-}=\frac{4n^{2}\pi^{2}}{d_{y}^{2}};$ (22)
here, the lowest-energy sub-band due to confinement along $y$ is $\pi^{2}/d_{y}^{2}$, with an energy gap of $3\pi^{2}/d_{y}^{2}$ to the next sub-band. Then, the matrix representation of the Hamiltonian in the chosen basis is given by $\hat{h}_{n,m}=\left\langle\Psi_{n}\left|\hat{\mathcal{H}}\right|\Psi_{m}\right\rangle$. It follows that
$\hat{h}_{n,m}$ $=\left\langle\psi_{n_{1}}\left|\hat{p}_{x}^{2}\right|\psi_{m_{1}}\right\rangle\delta_{n_{2},m_{2}}\hat{A}_{n_{3},m_{3}}+$
$\quad\delta_{n_{1},m_{1}}\delta_{n_{2},m_{2}}E_{y,m_{2}}\hat{B}_{n_{3},m_{3}}+$
$\quad\left\langle\psi_{n_{1}}\left|\hat{p}_{x}\right|\psi_{m_{1}}\right\rangle\left\langle\phi_{n_{2}}\left|\hat{p}_{y}\right|\phi_{m_{2}}\right\rangle\hat{C}_{n_{3},m_{3}}+$
$\quad\delta_{n_{1},m_{1}}\delta_{n_{2},m_{2}}\hat{D}_{n_{3},m_{3}}.$ (23)
It can be shown that the matrix element $\left\langle\phi_{n_{2}}\left|\hat{p}_{y}\right|\phi_{m_{2}}\right\rangle$ vanishes when $\phi_{n_{2}}$ and $\phi_{m_{2}}$ have the same parity (i.e., both even or both odd), and can be nonzero only when their parities differ. Therefore, $\hat{h}$ assumes the following block structure:
$\hat{h}=\left(\begin{array}[]{ccc}\hat{h}_{1D}^{\left(1\right)}&\hat{h}_{c}^{\left(1,2\right)}\\ \hat{h}_{c}^{\left(2,1\right)}&\hat{h}_{1D}^{\left(2\right)}\\ &&\ddots\end{array}\right).$ (24)
The $i$-th diagonal block corresponds to the one-dimensional effective Hamiltonian for the $i$-th sub-band, namely
$\hat{h}_{1D}^{\left(i\right)}=\hat{A}\hat{p}_{x}^{2}+\hat{B}E_{y,i}+\hat{D}.$ (25)
On the other hand, the off-diagonal block
$\hat{h}_{c}^{\left(i,j\right)}=\left\langle\phi_{i}\left|\hat{p}_{y}\right|\phi_{j}\right\rangle\hat{C}\hat{p}_{x}$ (26)
accounts for the coupling (in case of different parities) between the $i$-th and $j$-th sub-bands mediated by transverse mode mixing.
Crystal axis misalignment can result also in a $y$-dependent variation of the hopping amplitude at the interfaces. This physical scenario can be modeled by introducing a barrier potential acting at the left (L) and right (R) interfaces, expressed as
$f\left(x,y\right)=\hat{E}V_{L}\left(y\right)\delta\left(x+d_{x}/2\right)+\hat{F}V_{R}\left(y\right)\delta\left(x-d_{x}/2\right)$ (27)
with $\hat{E}=\left[\sum_{i}\nu_{L,i}\hat{L}_{i}^{2}+\tilde{g}_{L}\hat{\chi}_{z}\right]$, $\hat{F}=\left[\sum_{i}\nu_{R,i}\hat{L}_{i}^{2}+\tilde{g}_{R}\hat{\chi}_{z}\right]$, and the parameters $\nu$ and $\tilde{g}$ depending on the electronic and geometrical properties of the interfaces. Here, $V_{L}$ and $V_{R}$ are $y$-dependent barrier profiles that attain large amplitudes where the atomic separation suppresses the transmission probability. In this framework, using the $\phi_{n}$ representation for the transverse modes, we can rewrite the Hamiltonian in a way that each block (for the intra- and inter-band contributions) will include delta-function potentials along the $x$-direction depending on the character of the involved transverse modes. In particular, we have that the $i$-th diagonal block becomes
$\hat{h}_{1D}^{\left(i\right)}=$ $\hat{A}\hat{p}_{x}^{2}+\hat{B}E_{y,i}+\hat{D}+\hat{E}\tilde{V}_{L,i}\delta\left(x+d_{x}/2\right)+$
$\hat{F}\tilde{V}_{R,i}\delta\left(x-d_{x}/2\right),$ (28)
where $\tilde{V}_{L,i}=\left\langle\phi_{i}\left|V_{L}\left(y\right)\right|\phi_{i}\right\rangle$ ($\tilde{V}_{R,i}=\left\langle\phi_{i}\left|V_{R}\left(y\right)\right|\phi_{i}\right\rangle$) corresponds to the effective amplitude of a scattering delta potential present at the left (right) interface. At the same time, the off-diagonal block coupling the $i$-th and $j$-th mode takes the form
$\hat{h}_{c}^{\left(i,j\right)}=$ $\left\langle\phi_{i}\left|\hat{p}_{y}\right|\phi_{j}\right\rangle\hat{C}\hat{p}_{x}+\hat{E}\tilde{V}_{L,ij}\delta\left(x+d_{x}/2\right)+$
$\hat{F}\tilde{V}_{R,ij}\delta\left(x-d_{x}/2\right),$ (29)
where $\tilde{V}_{L,ij}=\left\langle\phi_{i}\left|V_{L}\left(y\right)\right|\phi_{j}\right\rangle$ and $\tilde{V}_{R,ij}=\left\langle\phi_{i}\left|V_{R}\left(y\right)\right|\phi_{j}\right\rangle$. At this stage, we exploit the hypothesis of ultra-thin limit regime by considering only the lowest-energy sub-band associated with confinement along $y$. This approximation is well-suited, as the energy separation between the ground and first excited sub-band scales as $d_{y}^{-2}$ and, therefore, is significantly large. It corresponds to set $n_{2}=m_{2}=1$, leading to an effective Hamiltonian for the system of the form
$\hat{h}_{eff}=$ $\hat{A}\hat{p}_{x}^{2}+\hat{B}E_{y,1}+\hat{D}+\hat{E}\tilde{V}_{L,1}\delta\left(x+d_{x}/2\right)+$
$\hat{F}\tilde{V}_{R,1}\delta\left(x-d_{x}/2\right).$ (30)
It is worth noting that, even in the limit of having only the lowest quantum well state occupied at the Fermi level with ideal interfaces ($\tilde{V}_{L,1}$=$\tilde{V}_{R,1}=0$), the effects of coupling between the confined states and the channel emerge through the term $\hat{B}E_{y,1}$. We also observe that each block-diagonal Hamiltonian shares identical symmetry-breaking terms linked to the $\hat{\chi}_{z}$ component. Then, the study of the effective 1D model Hamiltonian can capture the essential physics of the spin and orbital filtering effects in the transport of thin nanostructures. Consequently, the resulting effective Hamiltonian for the system is
{^$\begin{cases}\hat{\mathcal{H}}^{\text{I}}=\hat{p}_{x}^{2}\left(\sum_{i}a_{i}^{\text{I}}\hat{L}_{i}^{2}\right)+\sum_{i}\Delta_{i}^{\text{I}}\hat{L}_{i}^{2}\\ \hat{\mathcal{H}}^{\text{II}}=\hat{p}_{x}^{2}\left(\sum_{i}a_{i}^{\text{II}}\hat{L}_{i}^{2}+g_{z}\hat{\chi}_{z}\right)+\sum_{i}\Delta_{i}^{\text{II}}\hat{L}_{i}^{2}+\\ \,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,g_{0z}\hat{\chi}_{z}+\hat{V}\left(x\right)\\ \hat{\mathcal{H}}^{\text{III}}=\hat{p}_{x}^{2}\left(\sum_{i}a_{i}^{\text{III}}\hat{L}_{i}^{2}\right)+\sum_{i}\Delta_{i}^{\text{III}}\hat{L}_{i}^{2}\end{cases},$ (31)
where
@@
$\begin{cases}\hat{V}\left(x\right)=V_{1}\delta\left(x+d_{x}/2\right)\hat{E}+V_{2}\delta\left(x-d_{x}/2\right)\hat{F}\\ a_{i}^{\text{II}}=-a^{2}\left(\gamma_{i}^{\text{II}}\cos^{2}\theta+\delta_{i}^{\text{II}}\sin^{2}\theta\right)\\ g_{z}=-a^{2}\left(g_{1}\cos^{2}\theta+g_{2}\sin^{2}\theta\right)\\ \Delta_{i}^{\text{II}}=\alpha_{y}\gamma_{i}^{\text{II}}+\beta_{y}\delta_{i}^{\text{II}}+\xi_{i}^{\text{II}}\\ g_{0z}=\alpha_{y}g_{1}+\beta_{y}g_{2}\end{cases}$ (32)
with $\alpha_{y}=2-a^{2}E_{y,1}\sin^{2}\theta$ and $\beta_{y}=2-a^{2}E_{y,1}\cos^{2}\theta$. Further aspects related to this approach for the study of heterojunctions can be found in Refs. [42, 43].
@@
Hereafter, for clarity with respect to the notation in the main text, we assume that the electron propagates along the $z$-axis of a coordinate system $Oxyz$. To include the electronic spin degree, we employ the Pauli matrices: $\hat{S}_{i=x,y,z}=\frac{1}{2}\hat{\sigma}_{i=x,y,z}$. For the analysis of the transmission problem, the electron is prepared in a given state $\left|\Psi_{in}^{\text{I}}\right\rangle$ in region I. There, the Hamiltonian is expressed as ^$\hat{\mathcal{H}}^{\text{I}}\left(z\right)=\left[\hat{p}_{z}^{2}\left(\sum_{i}a_{i}^{\text{I}}\hat{L}_{i}^{2}\right)+\sum_{i}\Delta_{i}^{\text{I}}\hat{L}_{i}^{2}\right]\hat{\sigma}_{0}$, with $\hat{\sigma}_{0}=\hat{\sigma}_{i}^{2}$. The electron is thus injected from region I into region II, which spans from $z=-d/2$ to $z=d/2$. In this region, considering the factors involved in deriving the effective model for the mismatched heterostructure—taking into account various geometries or crystalline symmetries—the single-electron Hamiltonian will include all the components of the $\chi$-coupling, thus having the following expression:
@@
$\quad\left(\sum_{i}\Delta_{i}^{\text{II}}\hat{L}_{i}^{2}+\bm{g_{0}}\cdot\hat{\bm{\chi}}\right)\hat{\sigma}_{0}+\lambda\hat{\bm{L}}\cdot\hat{\bm{S}}+$
$\quad\left[V_{1}\delta\left(z+d/2\right)\hat{\mathcal{V}}_{1}+V_{2}\delta\left(z-d/2\right)\hat{\mathcal{V}}_{2}\right]\hat{\sigma}_{0}$ (33)
with $\bm{g}=\left(g_{x},g_{y},g_{z}\right)$, $\bm{g_{0}}=\left(g_{0x},g_{0y},g_{0z}\right)$, $\hat{\bm{\chi}}=\left(\left\{\hat{L}_{y},\hat{L}_{z}\right\},\left\{\hat{L}_{x},\hat{L}_{z}\right\},\left\{\hat{L}_{x},\hat{L}_{y}\right\}\right)$, $\lambda$ as the atomic spin-orbit coupling magnitude (in units of $\epsilon_{0}$), and $\hat{\mathcal{V}}_{i}=\sum_{j}\left(\nu_{i,j}\hat{L}_{j}^{2}+\tilde{g}_{i,j}\hat{\chi}_{j}\right)$. Finally, the electron can reach the last region of space, region III, where its orbital and spin angular momenta are measured, and the Hamiltonian can be expressed as ^$\hat{\mathcal{H}}^{\text{III}}\left(z\right)=\left[\hat{p}_{z}^{2}\left(\sum_{i}a_{i}^{\text{III}}\hat{L}_{i}^{2}\right)+\sum_{i}\Delta_{i}^{\text{III}}\hat{L}_{i}^{2}\right]\hat{\sigma}_{0}$. Thus, we can define the following overall Hamiltonian ^=^$\hat{\mathcal{H}}=\hat{\mathcal{H}}\left(z\right)$: